# New affine invariant ensemble samplers and their dimensional scaling

### v2

@@ Abstract.
We introduce new affine invariant ensemble samplers that are easy to construct and improve upon existing algorithms, especially for high-dimensional problems. Specifically, we propose a derivative-free ensemble side move sampler that performs favorably compared to popular samplers in the emcee package. Additionally, we develop a class of derivative-based ensemble Hamiltonian Monte Carlo (HMC) samplers with affine invariance, which outperform standard HMC without affine invariance when sampling highly skewed distributions. We provide asymptotic scaling analysis for high-dimensional Gaussian targets to further elucidate the properties of these affine invariant ensemble samplers. In particular, with derivative information, the affine invariant ensemble HMC can scale much better with dimension compared to derivative-free ensemble samplers.
@@ 1. Introduction
The concept of affine invariance [26] has played an important role in the development of efficient Markov chain Monte Carlo (MCMC) samplers as it provides robustness against anisotropy. Samplers are called affine invariant if they behave consistently across all coordinate systems related through affine transformations. In MCMC, affine invariance is typically achieved through ensemble samplers.
Affine invariant ensemble samplers have proven effective for routine Bayesian inference applications [20], but they may perform poorly in high-dimensional problems [32, 7]. This particularly applies to derivative-free samplers such as the stretch move sampler [26], widely used via the emcee package [20]. Indeed, compelling arguments suggest that “ensemble methods are doomed to fail in high dimensions” [7], based on insightful observations about typical sets: high-dimensional distributions (such as Gaussians) typically concentrate on thin shells, and the interpolation or extrapolation between two points in the stretch move—as well as in many other ensemble samplers—is unlikely to fall within this shell. As a result, small step sizes must be used, and the samplers effectively “devolve into random walks with poorly biased directional choices.”
In this paper, we introduce new affine invariant ensemble samplers and analyze their scaling behavior with dimension. We show that ensemble methods can, in fact, achieve the same high-dimensional scaling as their single-chain counterparts, for both derivative-free and derivative-based samplers. In particular, derivative-based ensemble samplers can break the random walk behavior and scale more favorably with dimension. Thanks to affine invariance, the hidden constants in the scaling are also insensitive to the problem’s condition number.
@@ 1.1.1. Derivative-free ensemble samplers
We first propose a derivative-free side move sampler that improves upon the stretch move by adopting a more favorable directional proposal in high dimensions. For Gaussian targets, we analyze the acceptance probability, which suggests that the optimal step size parameters (denoted $a-1$ for the stretch move and $\sigma$ for the side move) should scale as $d^{-1/2}$. This scaling matches that of the single-chain random walk [23, 55]. We further show that the expected squared jumping distance [43, 2, 48, 55] under the optimal step size is larger for the side move than for the stretch move. Our numerical experiments show that the autocorrelation times of both samplers scale linearly with $d$, and that the side move outperforms the stretch move by a factor of two or more across various $\sim$100-dimensional examples.
@@ 1.1.2. Derivative-based ensemble samplers
Beyond derivative-free samplers, it is well known that incorporating derivative information allows MCMC samplers to scale more favorably with dimension [47, 3]. In particular, the Hamiltonian Monte Carlo (HMC) sampler [41] can overcome random walk behavior and achieves state-of-the-art $d^{-1/4}$ scaling for the step size.
Whether it is possible to develop an efficient, affine invariant ensemble HMC sampler has remained an open question. We propose a new class of samplers in this vein, one of which achieves $d^{-1/4}$ scaling for step size on high-dimensional Gaussian targets. Importantly, our method employs a specific antisymmetric preconditioning strategy based on complementary ensembles to enable efficient parallelism — rather than relying on traditional mass matrices for preconditioning, as in standard HMC.
Our experiments demonstrate that affine invariant HMC samplers can outperform derivative-free samplers by an order of magnitude and achieve 10- to 100-fold reductions in autocorrelation time compared to standard HMC (without elaborate tuning) when sampling from $\sim$100-dimensional synthetic anisotropic distributions and distributions arising from stochastic PDEs, which are typically ill-conditioned.
Consequently, while ensemble samplers based on derivative-free interpolation or extrapolation inevitably devolve into random walks and may struggle in high dimensions, ensemble samplers that incorporate HMC’s derivative information avoid this limitation and are therefore not doomed to fail in high-dimensional settings.
@@ 1.2. Related work on affine invariant samplers
The concept of affine invariance, introduced to MCMC samplers by [26], draws inspiration from the empirical success of the Nelder-Mead simplex algorithm [42] in optimization. This idea has been extended to many other sampling and related areas, including data assimilation [44], annealed importance sampling [11], and variational inference [12], to list a few.
@@
The stretch move [26], which is related to the scaling-invariant $t$-walk move proposed in [15], remains one of the most widely used affine invariant ensemble samplers. Popularized through the emcee package [20], it has been extensively adopted in the astrophysics community. The stretch move is most effective in moderate dimensions (e.g., $d\leq 20$) but must be used cautiously for high-dimensional problems [32]. To address high-dimensional or infinite-dimensional problems, particularly in PDEs and inverse problems, researchers have developed hybrid approaches that combine function space MCMC [16] with ensemble samplers [17, 19].
@@
The stretch move is closely related to the snooker algorithm [24, 46], which uses ensembles to enable adaptive directional sampling. The ensemble covariance has also been used to design derivative-free affine invariant samplers, such as the walk move [26]. Another well-known ensemble sampler is differential evolution MCMC [6, 54], which is inspired by the differential evolution optimization algorithm [53]. The affine invariant side move we introduce in this paper is connected to both the walk move and differential evolution MCMC, as we detail in Section 3.
@@
Beyond derivative-free samplers, derivative-based affine invariant samplers have been developed, including those based on Riemannian geometry [25] and Newton-type directions [40, 50, 18, 14]. Affine invariance can also be achieved by combining ensemble covariance preconditioning with gradient-based directions [27, 35, 21, 22, 37, 12]. These methods establish mathematical links to gradient flows and ensemble Kalman filters via derivative-free approximations [21, 22, 12]. However, most of these approaches do not incorporate Metropolis adjustments and are therefore biased. Our focus in this paper is on adjusted, unbiased MCMC samplers.
@@
Among unbiased derivative-based MCMC samplers, HMC [41] achieves the state-of-the-art scaling with dimension [3]. However, HMC can struggle with anisotropic or ill-conditioned target distributions and requires adaptation of the mass matrix [8] to correct scale imbalances, as it is not affine invariant. Metrics based on Riemannian manifolds and Hessians have been used to make HMC affine invariant [25, 34, 33], but these approaches generally require costly second-order information. We believe that systematically developing affine invariant ensemble HMC methods — based solely on first-order information — represents a promising algorithmic opportunity to advance the state of ensemble samplers. We note that researchers have explored using ensemble chains to form a diagonal preconditioner based on individual variance [30], which leads to scale invariance — a property slightly weaker than affine invariance. Related work has also focused on tuning-free multichain adaptive HMC methods that utilize parallel hardware [28, 52, 51].
@@ 2. Affine Invariance
We review the mathematical concept of affine invariance. Following the seminal work of Goodman and Weare [26], we express a general MCMC sampler as
@@ 3.1. Basic side move
where $\sigma$ is a user-specified scalar parameter, and $\xi\sim\mathcal{N}(0,1)$ is drawn from a normal distribution. We will discuss connections to walk move [26] and differential evolution [53, 6, 54] in Sections 3.3 and 3.4. In Section 3.5, we study the high dimensional scaling behavior, which suggests a choice of $\sigma=1.687d^{-1/2}$.
@@ 3.2. Parallel side move sampler
The side move algorithm can be efficiently parallelized. However, this needs to be done carefully to avoid violating detailed balance. Our approach is similar to that used in the emcee package [20]. Specifically, we divide the ensemble into two groups:
@@ 3.3. Comparison to other affine invariant moves
One of the most popular affine invariant sampler is based on stretch move [26]. In the basic stretch move, at step $m$, one proposes
@@
where $a$ is the stretch parameter with a recommended default value $a=2$ (see [20]). To maintain detailed balance, the proposal is accepted with probability
@@
We note that other affine invariant moves based on empirical covariance exist, such as the walk move [26]. In the walk move, we select a subset $S$ of particles which are different from $\mathbf{x}_{i}(m)$. Denote the empirical mean by $\mathbf{m}_{S}$. The proposal is
@@
The empirical covariance accounts for global features of the distribution if $|S|$ is large. As discussed in [35, 45], local features could be more representative for certain distributions, which may explain the wider popularity of the stretch move in the literature [20]. We note that when $|S|=2$, the walk move can be shown to be equivalent to the side move with a specific step size. This can be demonstrated by noting:
@@ 3.4. Connection to differential evolution
Affine invariant ensemble MCMC [26] was motivated by optimization algorithms. The side move is also connected to methods in the optimization literature, specifically the differential evolution algorithm [53], where particle differences are used to guide the exploration of the function. Differential evolution MCMC [6, 54] has been developed and has found numerous successful applications. In detail, differential evolution MCMC employs the proposal
@@ 3.5. Analysis of high dimensional scaling
We study the high dimensional scaling of the algorithm for Gaussians at the stationary phase. Since the algorithm is affine invariant, we consider isotropic Gaussians without loss of generality. The analysis suggests $\sigma$ (in side move) and $a-1$ (in stretch move) to scale with $d^{-1/2}$ in high dimensions. We also study the limit of expected squared jumped distance [43, 2, 48, 55], which has been used to derive optimal parameters of MCMC. The proof of this proposition can be found in Appendix B.
@@
The scaling of $\sigma\sim d^{-1/2}$ for the side move is natural since it resembles a basic random walk, aligning with existing scaling results for non-affine-invariant random walk Metropolis [23, 55]. The above result applies to arbitrary Gaussians given the affine invariance property of the side move. The scaling $a-1\sim d^{-1/2}$ for the stretch move is perhaps less obvious. The presence of the $Z^{d-1}$ factor is the key element leading to such scaling.
@@ 4. Derivative-based Affine Invariant HMC Samplers
The samplers in the previous section are derivative-free, making them simple, convenient, and widely used in applications. However, when derivative information is available, derivative-based samplers such as Langevin [49] and Hamiltonian Monte Carlo (HMC) [41] have been shown to scale better with dimension [47, 3].
There has been considerable interest in developing affine invariant Langevin samplers based on Riemannian geometry and empirical covariance preconditioning [25, 27, 35, 21, 22]. To the best of our knowledge, it remains unknown whether an adjusted ensemble HMC that is affine invariant can be efficiently developed. Since HMC is one of the state-of-the-art samplers that overcomes random walk-like behavior and scales favorably with dimension, developing an affine invariant ensemble HMC is important for further understanding and advancement of affine invariant samplers. This section aims to discuss the design principles and propose new samplers of this kind.
@@ 4.1. Hamiltonian Monte Carlo
The key to the derivation of HMC is the property that $\mathsf{P}\mathsf{L}_{h}^{n}=(\mathsf{P}\mathsf{L}_{h}^{n})^{-1}$ so the detailed balance for the second step is valid. Modern development of HMC focuses on tuning the stepsize and integration time [29, 8], among many others.
@@ 4.2. Covariance preconditioning and its challenge
We note that when marginalized over all $\mathbf{p}_{i}$, the above joint distribution does not give the correct marginal distribution for $\mathbf{x}_{i}$ due to the existence of normalization constants. In fact, we need to additionally add $-\frac{N}{4}\log\det\mathrm{Cov}_{S^{(0)}}-\frac{N}{4}\log\det\mathrm{Cov}_{S^{(1)}}$ to the potential to address this issue. The modified final joint distribution is
@@
(4.4) $V_{\star}(\mathbf{x}_{1},...,\mathbf{x}_{N},\mathbf{p}_{1},...\mathbf{p}_{N})=$ $\sum_{i=1}^{N/2}\left(V(\mathbf{x}_{i})+\frac{1}{2}\mathbf{p}_{i}^{T}\mathrm{Cov}_{S^{(1)}}\mathbf{p}_{i}\right)$
@@
  • We fix particles in $S^{(1)}$, and update particles in $S^{(0)}$: for each $21\leq i\leq N/$ 2, we sample $\mathbf{p}_{i}\sim\mathcal{N}(0,\mathrm{Cov}_{S^{(1)}}^{-1})$, run leapfrog approximation of the dynamics
@@
(4.5) $(\mathbf{x}_{1},...,\mathbf{x}_{N})\to(\mathbf{y}_{1},...,\mathbf{y}_{N})=(A\mathbf{x}_{1}+\mathbf{b},...,A\mathbf{x}_{N}+\mathbf{b})$
$(\mathbf{p}_{1},...,\mathbf{p}_{N})\to(\mathbf{r}_{1},...,\mathbf{r}_{N})=(A^{-T}\mathbf{p}_{1},...,A^{-T}\mathbf{p}_{N})\,.$
By the change of variables, we get (for $21\leq i\leq N/$ 2)

### v3

@@ Abstract.
We introduce new affine invariant ensemble Markov chain Monte Carlo (MCMC) samplers that are easy to construct and improve upon existing methods, especially for high-dimensional problems. We first propose a simple derivative-free side move sampler that improves upon popular samplers in the emcee package by generating more effective proposal directions. We then develop a class of derivative-based affine invariant ensemble Hamiltonian Monte Carlo (HMC) samplers based on antisymmetric preconditioning using complementary ensembles, which outperform standard, non-affine-invariant HMC when sampling highly anisotropic distributions. We provide asymptotic scaling analysis for high-dimensional Gaussian targets to further elucidate the properties of these affine invariant ensemble samplers. In particular, with derivative information, the affine invariant ensemble HMC can scale much better with dimension compared to derivative-free ensemble samplers.
@@ 1. Introduction
The concept of affine invariance [goodman2010ensemble] has played an important role in the development of efficient Markov chain Monte Carlo (MCMC) samplers as it provides robustness against anisotropy and obviates the need for problem-specific preconditioning. Samplers are called affine invariant if they behave consistently across all coordinate systems related through affine transformations. In practice, affine invariance in MCMC is commonly achieved through ensemble samplers, which propose updates based on relative configurations of multiple interacting chains.
Affine invariant ensemble samplers have proven effective for routine Bayesian inference applications [foreman2013emcee], but they may perform poorly in high-dimensional problems [huijser2015properties, carpenter2017ensemble]. This particularly applies to derivative-free samplers such as the stretch move sampler [goodman2010ensemble], widely used via the emcee package [foreman2013emcee]. Indeed, compelling arguments suggest that “ensemble methods are doomed to fail in high dimensions” [carpenter2017ensemble], based on insightful observations about typical sets: high-dimensional distributions such as Gaussians typically concentrate on thin shells, and the interpolation or extrapolation between points in the stretch move—as well as in many other ensemble samplers—is unlikely to fall within this shell. As a result, small step sizes are required to maintain reasonable acceptance rates, and the samplers effectively “devolve into random walks with poorly biased directional choices.”
This paper revisits these concerns and shows that the situation is more nuanced. We introduce new affine invariant ensemble samplers and analyze their scaling behavior with dimension. We demonstrate that ensemble methods can, in fact, achieve the same high-dimensional scaling as their single-chain counterparts, both in the derivative-free and derivative-based cases and at the stationary phase. In particular, ensemble samplers that incorporate gradient information can overcome random-walk behavior and scale more favorably with dimension. Moreover, due to affine invariance, the constants hidden in the asymptotic scaling are insensitive to the condition number of the target distribution.
@@ 1.1.1. Derivative-free ensemble samplers
We first propose a derivative-free side move ensemble sampler that improves upon the stretch move by adopting a more favorable proposal direction in high dimensions. For Gaussian targets, we analyze the acceptance probability and show that the optimal step-size parameters—denoted a1a-1 for the stretch move and $\sigma$ for the side move—scale as $d^{-1/2}$. This scaling matches that of the single-chain random walk Metropolis algorithm [gelman1997weak, yang2020optimal]. We further show that, under optimal scaling, the expected squared jumping distance [pasarica2010adaptively, atchade2011towards, roberts2014minimising, yang2020optimal] is larger for the side move than for the stretch move.
Our numerical experiments confirm these theoretical predictions: the integrated autocorrelation times of both samplers scale linearly with the dimension $d$, while the side move consistently outperforms the stretch move—often by a factor of two or more—across a range of problems in dimensions on the order of $100$.
@@ 1.1.2. Derivative-based ensemble samplers
Beyond derivative-free methods, it is well known that incorporating gradient information allows MCMC samplers to scale more favorably with dimension [roberts1998optimal, beskos2013optimal]. In particular, Hamiltonian Monte Carlo (HMC) [neal2011mcmc] avoids random-walk behavior and achieves the state-of-the-art $d^{-1/4}$ scaling for the step size.
Whether one can construct an efficient affine invariant ensemble HMC sampler has remained an open question. In this work, we propose a new class of affine invariant ensemble HMC methods, one of which attains $d^{-1/4}$ step-size scaling for high-dimensional Gaussian targets. The key idea is to employ an antisymmetric preconditioning strategy derived from complementary ensembles, enabling efficient parallel updates without relying on traditional mass-matrix tuning as in standard HMC.
Our experiments demonstrate that these affine invariant HMC samplers can outperform derivative-free ensemble methods by an order of magnitude and achieve 10- to 100-fold reductions in autocorrelation time compared to standard HMC (without elaborate tuning of mass matrices) when sampling from $\sim$100-dimensional synthetic anisotropic distributions and from distributions arising in stochastic PDE models, which are typically ill-conditioned.
In summary, while ensemble samplers based purely on derivative-free interpolation or extrapolation inevitably exhibit random-walk behavior and struggle in high dimensions, ensemble samplers that incorporate derivative information in the spirit of HMC avoid this limitation and are therefore not doomed to fail in high-dimensional settings.
@@ 1.2. Related work on affine invariant samplers
The concept of affine invariance, introduced to MCMC samplers by [goodman2010ensemble], draws inspiration from the empirical success of the Nelder-Mead simplex algorithm [nelder1965simplex] in optimization. This idea has been extended to many other sampling and related areas, including data assimilation [reich2015probabilistic], annealed importance sampling [chen2024ensemble], and variational inference [chen2023sampling], to list a few.
@@
The stretch move [goodman2010ensemble], which is related to the scaling-invariant $t$-walk move proposed in [christen2007general], remains one of the most widely used affine invariant ensemble samplers. Popularized through the emcee package [foreman2013emcee], it has been extensively adopted in the astrophysics community. The stretch move is most effective in moderate dimensions (e.g., $d\leq 20$) but must be used cautiously for high-dimensional problems [huijser2015properties]. To address high-dimensional or infinite-dimensional problems, particularly in PDEs and inverse problems, researchers have developed hybrid approaches that combine function space MCMC [cotter2013mcmc] with ensemble samplers [coullon2021ensemble, dunlop2022gradient].
@@
The stretch move is closely related to the snooker algorithm [gilks1994adaptive, roberts1994convergence], which uses ensembles to enable adaptive directional sampling. The ensemble covariance has also been used to design derivative-free affine invariant samplers, such as the walk move [goodman2010ensemble]. Another well-known ensemble sampler is differential evolution MCMC [braak2006markov, vrugt2009accelerating], which is inspired by the differential evolution optimization algorithm [storn1997differential]. The affine invariant side move we introduce in this paper is connected to both the walk move and differential evolution MCMC, as we detail in Section 3.
@@
Beyond derivative-free samplers, derivative-based affine invariant samplers have been developed, including those based on Riemannian geometry [girolami2011riemann] and Newton-type directions [martin2012stochastic, simsekli2016stochastic, detommaso2018stein, chewi2020exponential]. Affine invariance can also be achieved by combining ensemble covariance preconditioning with gradient-based directions [greengard2015ensemblized, leimkuhler2018ensemble, garbuno2020interacting, garbuno2020affine, liu2022second, chen2023sampling]. These methods establish mathematical links to gradient flows and ensemble Kalman filters via derivative-free approximations [garbuno2020interacting, garbuno2020affine, chen2023sampling]. However, most of these approaches do not incorporate Metropolis adjustments and are therefore biased. Our focus in this paper is on adjusted, unbiased MCMC samplers.
@@
Among unbiased derivative-based MCMC samplers, HMC [neal2011mcmc] achieves the state-of-the-art scaling with dimension [beskos2013optimal]. However, HMC can struggle with anisotropic or ill-conditioned target distributions and requires careful adaptation of the mass matrix [carpenter2017stan] to correct scale imbalances, as it is not affine invariant. Metrics based on Riemannian manifolds and Hessians have been used to make HMC affine invariant [girolami2011riemann, lee2018convergence, kook2022sampling], but these approaches generally require costly second-order information. We believe that systematically developing affine invariant ensemble HMC methods — based solely on first-order information — represents a promising algorithmic opportunity to advance the state of ensemble samplers. Related work has explored diagonal preconditioning based on ensemble variance [hoffman2022tuning], yielding scale invariance—a weaker notion than affine invariance—as well as tuning-free multichain adaptive HMC methods designed for parallel architectures [hoffman2021adaptive, sountsov2021focusing, sountsov2024running].
@@ 2. Affine Invariance
We review the mathematical concept of affine invariance. Following the seminal work of Goodman and Weare [goodman2010ensemble], we express a general MCMC sampler as
@@ 3.1. Basic side move
where $\sigma$ is a user-specified scalar parameter, and $\xi\sim\mathcal{N}(0,1)$ is drawn from a normal distribution. We will discuss connections to walk move [goodman2010ensemble] and differential evolution [storn1997differential, braak2006markov, vrugt2009accelerating] in Sections 3.3 and 3.4. In Section 3.5, we study the high dimensional scaling behavior, which suggests a choice of $\sigma=1.687d^{-1/2}$.
@@ 3.2. Parallel side move sampler
The side move algorithm can be efficiently parallelized. However, this needs to be done carefully to avoid violating detailed balance. Our approach is similar to that used in the emcee package [foreman2013emcee]. Specifically, we divide the ensemble into two groups:
@@ 3.3. Comparison to other affine invariant moves
One of the most popular affine invariant sampler is based on stretch move [goodman2010ensemble]. In the basic stretch move, at step $m$, one proposes
@@
where $a$ is the stretch parameter with a recommended default value $a=2$ (see [foreman2013emcee]). To maintain detailed balance, the proposal is accepted with probability
@@
We note that other affine invariant moves based on empirical covariance exist, such as the walk move [goodman2010ensemble]. In the walk move, we select a subset $S$ of particles which are different from $\mathbf{x}_{i}(m)$. Denote the empirical mean by $\mathbf{m}_{S}$. The proposal is
@@
The empirical covariance accounts for global features of the distribution if $|S|$ is large. As discussed in [leimkuhler2018ensemble, reich2021fokker], local features could be more representative for certain distributions, which may explain the wider popularity of the stretch move in the literature [foreman2013emcee]. We note that when $|S|=2$, the walk move can be shown to be equivalent to the side move with a specific step size. This can be demonstrated by noting:
@@ 3.4. Connection to differential evolution
Affine invariant ensemble MCMC [goodman2010ensemble] was motivated by optimization algorithms. The side move is also connected to methods in the optimization literature, specifically the differential evolution algorithm [storn1997differential], where particle differences are used to guide the exploration of the function. Differential evolution MCMC [braak2006markov, vrugt2009accelerating] has been developed and has found numerous successful applications. In detail, differential evolution MCMC employs the proposal
@@ 3.5. Analysis of high dimensional scaling
We study the high dimensional scaling of the algorithm for Gaussians at the stationary phase. Since the algorithm is affine invariant, we consider isotropic Gaussians without loss of generality. The analysis suggests $\sigma$ (in side move) and a1a-1 (in stretch move) to scale with $d^{-1/2}$ in high dimensions. We also study the limit of expected squared jumped distance [pasarica2010adaptively, atchade2011towards, roberts2014minimising, yang2020optimal], which has been used to derive optimal parameters of MCMC. The proof of this proposition can be found in Appendix B.
@@
The scaling of $\sigma\sim d^{-1/2}$ for the side move is natural since it resembles a basic random walk, aligning with existing scaling results for non-affine-invariant random walk Metropolis [gelman1997weak, yang2020optimal]. The above result applies to arbitrary Gaussians given the affine invariance property of the side move. The scaling $a-1\sim d^{-1/2}$ for the stretch move is perhaps less obvious. The presence of the $Z^{d-1}$ factor is the key element leading to such scaling.
@@ 4. Derivative-based Affine Invariant HMC Samplers
The samplers in the previous section are derivative-free, making them simple, convenient, and widely used in applications. However, when derivative information is available, derivative-based samplers such as Langevin [roberts1996exponential] and Hamiltonian Monte Carlo (HMC) [neal2011mcmc] have been shown to scale better with dimension [roberts1998optimal, beskos2013optimal].
There has been considerable interest in developing affine invariant Langevin samplers based on Riemannian geometry and empirical covariance preconditioning [girolami2011riemann, greengard2015ensemblized, leimkuhler2018ensemble, garbuno2020interacting, garbuno2020affine]. To the best of our knowledge, it remains unknown whether an adjusted ensemble HMC that is affine invariant can be efficiently developed. Since HMC is one of the state-of-the-art samplers that overcomes random walk-like behavior and scales favorably with dimension, developing an affine invariant ensemble HMC is important for further understanding and advancement of affine invariant samplers. This section aims to discuss the design principles and propose new samplers of this kind.
@@ 4.1. Hamiltonian Monte Carlo
The key to the derivation of HMC is the property that $\mathsf{P}\mathsf{L}_{h}^{n}=(\mathsf{P}\mathsf{L}_{h}^{n})^{-1}$ so the detailed balance for the second step is valid. Modern development of HMC focuses on tuning the stepsize and integration time [hoffman2014no, carpenter2017stan], among many others.
@@ 4.2. Covariance preconditioning and its challenge
We note that when marginalized over all $\mathbf{p}_{i}$, the above joint distribution does not give the correct marginal distribution for $\mathbf{x}_{i}$ due to the existence of normalization constants. In fact, we need to additionally add $\frac{N}{4}\log\det\mathrm{Cov}_{S^{(0)}}-\frac{N}{4}\log\det\mathrm{Cov}_{S^{(1)}}$ to the potential to address this issue. The modified final joint distribution is
@@
(4.4) $V_{\star}(\mathbf{x}_{1},.,\mathbf{x}_{N},\mathbf{p}_{1},.\mathbf{p}_{N})=$ $\sum_{i=1}^{N/2}\left(V(\mathbf{x}_{i})+\frac{1}{2}\mathbf{p}_{i}^{T}\mathrm{Cov}_{S^{(1)}}\mathbf{p}_{i}\right)$
@@
  • We fix particles in $S^{(1)}$, and update particles in $S^{(0)}$: for each $1\leq i\leq N/2$, we sample $\mathbf{p}_{i}\sim\mathcal{N}(0,\mathrm{Cov}_{S^{(1)}}^{-1})$, run leapfrog approximation of the dynamics
@@
(4.5) $(\mathbf{x}_{1},.,\mathbf{x}_{N})\to(\mathbf{y}_{1},.,\mathbf{y}_{N})=(A\mathbf{x}_{1}+\mathbf{b},.,A\mathbf{x}_{N}+\mathbf{b})$
$(\mathbf{p}_{1},.,\mathbf{p}_{N})\to(\mathbf{r}_{1},.,\mathbf{r}_{N})=(A^{-T}\mathbf{p}_{1},.,A^{-T}\mathbf{p}_{N})\,.$
By the change of variables, we get (for $1\leq i\leq N/2$)